### v1

# A Scoping Review of Synthetic Data Generation for Biomedical Research and Applications

Abstract
Synthetic data generation––mitigating data scarcity, privacy concerns, and data quality challenges in biomedical fields––has been facilitated by rapid advances of large language models (LLMs). This scoping review follows PRISMA-ScR guidelines and synthesizes 59 studies, published between 2020 and 2025 and collected from PubMed, ACM, Web of Science, and Google Scholar. The review systematically examines biomedical research and application trends in synthetic data generation, emphasizing clinical applications, methodologies, and evaluations. Our analysis identifies data modalities of unstructured texts (78.0%), tabular data (13.6%), and multimodal sources (8.4%); generation methods of prompting (72.9%), fine-tuning (22.0%) LLMs and specialized model (5.1%); and heterogeneous evaluations of intrinsic metrics (27.1%), human-in-the-loop assessments (55.9%), and LLM-based evaluations (13.6%). The analysis addresses current limitations in what, where, and how health professionals can leverage synthetic data generation for biomedical domains. Our review also highlights challenges in adaption across clinical domains, resource and model accessibility, and evaluation standardizations.
Introduction
Data is the key to train reliable AI models for broad biomedical and clinical applications, such as medical diagnosis [1, 2], therapeutic treatments [3], and drug discovery [4]. However, obtaining massive, privacy-preserving, and high-quality data faces critical challenges, such as data availability, noisy and missing data, and legal regulations. Increasing biomedical studies are turning attentions to synthetic data generation, a process of creating artificial datasets that accurately replicate the statistical and structural properties of real-world data. Nonetheless, creating high-quality synthetic data remains challenging due to the inherent heterogeneity, complexity, and variability characteristic of biomedical data. Large language models (LLMs), as a generative AI method, offer a promising solution by enhancing the accuracy and diversity of synthetic datasets through advanced learning algorithms.
Synthetic data generation in the biomedical and clinical fields has achieved significant advances by LLMs (e.g., GPT-4 [5] and Llama 3 [6]) in the recent years, as shown in Figure 2. For example, LLMs have been applied successfully to obtain clinical narratives and simulate real patient records of mental health for diagnosis and behavior analysis [7], such as predicting suicide [8] and depression patterns [9]. The strong generation capability not only relieve data privacy concerns [10] to facilitate model developments but also may promote data quality [11] to support clinical decision-making. For example, synthetic text corpora have augmented classification models for cardiovascular and Alzheimer’s disease diagnosis [12]; A study develops a multi-agent dialogue generation framework (NoteChat) that creates synthetic patient-physician conversations to improve clinical documentation [13]. Nevertheless, questions and uncertainties remain regarding the best practices, validation approaches, and specific application areas for those LLM-generated synthetic data.
The goal of this scoping review is to comprehensively summarize and assess recent research publications on the biomedical and clinical applications leveraging LLMs for synthetic data generation. While multiple reviewing studies [14, 15, 16] have covered the LLM-generated synthetic data for general domains, studies on biomedical and clinical domains are still not available. For example, a close study [16] examines recent developments of generation algorithms and models on news and social media domains instead of biomedical fields. Specifically, through this study, we will identify and present the current state-of-the-art models, technical methods, evaluations or assessments of synthetic data quality, and gaps and opportunities for future research. We seek to answer a concrete yet unsolved question: what biomedical and clinical applications can be effectively addressed using LLMs-generated synthetic data, and by how?
Results
Our search and initial review resulted in 132 articles and finalized with 59 articles for this scoping review. XH, HW, and HR reviewed full texts of the articles and removed 73 articles due to ineligible article types (n=2), no large language model (n=7), unrelated to biomedical data generation (n=18), and not peer-reviewed and published in a journal or conference proceedings (n=46). Fifty-nine articles met our criteria and were included in our subsequent analyses. In terms of publication venues and formats, the included literature consists of 43 journal articles (72.9%) and 16 peer-reviewed conference papers (27.1%), leaving 59 studies to be included in this scoping review. Figure 1 shows that the study number of biomedical synthetic data generation climb steadily over years. Particularly, the substantial growth after 2022 reflects a particularly evolving trend of synthetic biomedical data generation via LLM, an emerging generation tool in recent years.
Figure 1: Publications between Jan 01, 2020 and April 05, 2025. Orange and blue colors refer to peer-reviewed conference and journal publications, respectively. We can observe a surging increase on the biomedical synthetic data related studies.
We present an overview of recent progress in synthetic biomedical data generation and application from three perspectives, aiming to provide readers with an insightful understanding of current practices and inform future research directions:
  1. 1.
    Synthetic data types and medical applications: Research has explored broad synthetic data types of clinical text, tabular data, and multimodal information (e.g., images, audio, and sensor signals). These data types cover a range of applications, including EHR analysis, medical imaging, telemedicine, mental health, and clinical trial matching, addressing challenges related to data scarcity and privacy restrictions.
  2. 2.
    Synthetic data generation methodology: Evolving methodologies encompass prompting-based generation, knowledge infusion, and multi-agent and multimodal approaches. These advancements enhance the semantic consistency, efficiency, and scalability of synthetic biomedical data generation.
  3. 3.
    Quality assessments and evaluation metrics: Ensuring data quality and utility remains a priority. Current studies employ evaluation strategies targeting fidelity, utility, and privacy protection. Such strategies integrate statistical metrics, human-in-the-loop review, automated model assessments, and privacy testing to establish a robust framework for verifying the reliability and compliance of synthetic data.
Table 1: A summary of data types, generation Methods, and medical applications for our collected studies. We use language codes for each study’s data language, including English (EN), Dutch (NL), French (FR), Chinese (ZH), and Arabic (AR). The K (1,000) shortens spaces of the size column. We use acronyms for medical applications, including QA (question answering), IE (information extraction), and SDoH (social determinants of health). We use the “-” to indicate not specified information.
Study Modality Lang. Architecture Approach Size (K) Year Accessible Purpose Medical Application
[17] Text EN Transformer Fine-tuning 500 2020 No Training Biomedical QA
[18] Text EN Transformer Specialized model 11 2020 Yes Training Phenotype inference
[19] Text EN GPT Fine-tuning <1K 2021 Yes Training Clinical IE
[20] Text NL GPT, RNN Fine-tuning 355 2021 Yes Training De-identification
[21] Text EN GPT Prompting 45 2021 No Training Dialogue summarization
[22] Text EN GPT Fine-tuning 48 2022 No Training Readmission prediction
[23] Text EN GPT Fine-tuning 200 2023 Yes Training General
[24] Text FR GPT Fine-tuning 5 2023 Yes Training Clinical IE
[25] Text EN GPT Fine-tuning 5 2023 No Training Syndromic detection
[26] Text EN GPT Fine-tuning <1K 2023 Yes Training Assisted suicide
[27] Tabular EN Transformer Specialized model 976 2023 Yes Training General
[12] Text EN GPT Prompting 32 2023 No Training Alzheimer’s disease
[28] Text EN GPT Prompting <1K 2024 Yes Training Discharge summary
[29] Text EN GPT Prompting 2 2024 Yes Training PTSD diagnosis
[30] Image ZH, EN GPT Prompting 4,000 2024 No Training Medical imaging
[31] Text EN GPT Prompting 9 2024 Yes Training Acute renal failure
[10] Text EN GPT Prompting 4 2024 No Privacy Phenotype inference
[32] Text EN GPT Prompting 2 2024 No Training Cohort selection
[33] Text EN GPT Prompting <1K 2024 No Supplement Cohort selection
[34] Text DE GPT Prompting 4 2024 No Training Emergency medicine
[35] Text EN GPT Prompting 19 2024 Yes Training Psychological therapy
[36] Text EN GPT Fine-tuning 30 2024 Yes Training Medical QA
[37] Text EN Transformer Fine-tuning 2 2024 Yes Training Medical QA
[11] Text EN GPT Prompting 5 2024 Yes Training General
[38] Text EN GPT Prompting 2 2024 Yes Supplement SDoH extraction
[39] Tabular EN GPT Prompting <1K 2024 No Supplement Medical education
[13] Text EN GPT Prompting 30 2024 Yes Supplement Dialogue generation
[40] Text EN GPT Prompting 14 2024 No Training SDoH extraction
[41] Tabular EN GPT Prompting - 2024 No Training Gait rehabilitation
[42] Text EN GPT Prompting 158 2024 Yes Training General
[43] Audio EN GPT Prompting 87 2024 Yes Supplement EMS assistance
[44] Text EN GPT Prompting 37 2024 Yes Training Clinical IE
[8] Text EN GPT, T5 Fine-tuning 148 2024 No Supplement Suicide prediction
[45] Text EN GPT Prompting <1K 2024 Yes Training Cancer symptom extraction
[46] Text EN GPT Prompting 5 2024 Yes Training Clinical summary
[47] Text EN GPT Prompting 5 2024 Yes Training Medical QA
[48] Tabular - GPT Prompting <1K 2024 No Supplement Nursing care
[49] Text EN GPT Prompting <1K 2024 Yes Training Patient portal triage
[50] Text EN GPT Prompting 100 2024 No Training Report generation
[51] Text AR GPT Prompting - 2024 Yes Supplement Health chatbot
[52] Tabular EN GPT Prompting <1K 2024 No Supplement Clinical trial simulation
[9] Text EN GPT Prompting 3 2024 No Supplement Depression symptom
[53] Text EN Transformer Prompting 4 2024 No Supplement Autism diagnosis
[54] Text EN GPT Prompting <1K 2025 No Privacy Stroke prediction
[55] Audio EN GPT Prompting 3 2025 No Training Dialogue analysis
[56] Text EN GPT Prompting 459 2025 No Supplement Phenotype ontology
[57] Text EN GPT Prompting 5 2025 No Training Radiology prediction
[58] Text ZH GPT Prompting 270 2025 No Training Anesthesia QA
[59] Tabular - Transformer Specialized model - 2025 Yes Supplement Patient outcome prediction
[60] Text ZH GPT Prompting 6 2025 Yes Training Mental health IE
[61] Tabular EN GPT Prompting 6 2025 Yes Privacy Data scarcity
[62] Text EN GPT Prompting 88 2025 No Privacy Phenotype inference
[63] Text EN GPT Prompting 5 2025 Yes Training Medical reasoning
[64] Text ZH GPT Prompting 45 2025 No Training Database search
[65] Text EN GPT Prompting 74 2025 No Training Metastases detection
[66] Image EN Transformer Fine-tuning 2 2025 Yes Supplement Colonoscopy analysis
[67] Image EN GPT Prompting - 2025 No Supplement Pathology analysis
[68] Text ET GPT Fine-tuning 4 2025 No Training Clinical IE
[69] Tabular EN GPT Prompting 10 2025 No Training Patient outcome prediction
Synthetic Data Types and Medical Applications
Increasing model developments for biomedical applications lead to the surging needs of synthetic data [18, 20]. In this section, we provide an overview of recent studies using synthetic data and address three critical questions: 1) What types of synthetic data are being generated? 2) What modalities and clinical tasks do these datasets target? 3) Are those data resources accessible for reusable research, such as benchmarking and model training? To answer those questions, we examine several aspects of the collected studies and summarize them in Table 1, including data modality, data language, data size, published year, medical application, and data or model accessibility. We aim to provide insights into current and future trends of data types and biomedical applications.
Biomedical synthetic data has shown diverse medical topics, languages, and modalities. Most of the studies (55.9%, $n=33$) are for general topics, while others include specific medical areas, such as suicide [26, 8], colorectal cancer [66], and radiology [57]. For example, studies generate synthetic data for augmenting diagnosis accuracy for Post-traumatic stress disorder (PTSD) [29] and Autism Spectrum Disorders (ASD) [53]. English is the predominant language (84.7%, $n=50$), though there are also studies in French [24], Chinese [30, 58, 60, 64], German [34] and Arabic [51], reflecting some progress toward non-English-speaking patients. the study [34] deployed multilingual LLMs to generate German data to simulate realistic emergency-medical dialogues between ambulance staff and patients. Most datasets focus on unstructured text modality (e.g., clinical narratives [13, 21, 50] and discharge summaries [28]), while a smaller but growing subset includes tabular [27, 39, 41, 48, 52, 59, 61, 69], image [30, 66, 67], and audio [55]. For example, Theodorou et al. [27] generate novel tabular datasets of high-dimensional longitudinal EHR records to provide realistic, privacy-preserving alternatives for machine learning and statistical analysis, while Ejiga et al. [66] generate synthetic colonoscopy image datasets via fine-tuned text-to-image synthesis to augment training data for robust colorectal cancer detection and precise polyp segmentation. Data sizes vary across health issues (e.g., cancer and mental health), yet the majority has relatively small sets with fewer than 10K samples, indicating the critical utility of synthetic data under low-resourced scenarios. For example, FairPlay [59] generates synthetic data from authentic EHR records to bolster under-represented patient subgroups, boosting mortality-prediction F1 scores by up to 21% and markedly narrowing performance gaps across different demographic groups. The various synthetic datasets highlight a growing community to deploy the biomedical synthetic data.
Synthetic data generation in biomedical research increasingly supports a wide array of clinical applications. Those tasks cover diverse medical needs, including phenotype classification [45, 53, 25, 12, 65], PSTD symptom extraction [29], de-identification (SynNote) [20], clinical note summarization  [28, 46], and mental health assessment  [9, 35, 8]. For example, CALLM [29] constructs clinical interview datasets to aid PTSD diagnosis by generating synthetic transcripts; and Barabadi et al. [65] generated synthetic radiology reports to enable automatic detection of metastatic sites in cancer patients. Those applications broadening the utility of synthetic data for health challenges, such as data shortage [29] and privacy concerns [20, 10, 62, 61, 54]. However, accessibility remains a central concern for advancing synthetic data utility. Table 1 shows 50.8% ($n=30$) of the studies are explicitly described as accessible, providing open or partially open resources for reproducible development, while others’ accessibility details or licensing remain unclear. Open datasets such as Syn-HPI [19], ClinGen [11], and Syn-EMS-Audio [43] serve as synthetic data benchmarks for evaluating new models and methods, while others—particularly those derived from sensitive clinical domains—may be restricted or anonymized to protect patient privacy. Ensuring clear, well-documented access protocols and licenses will be essential for fostering broader collaboration and translational impacts of synthetic data in biomedical research.
Synthetic Data Generation Methodology
Emerging Transformer-based large language models (LLMs) have significantly advanced synthetic data generation methodologies in biomedical research, enabling diverse applications, as summarized in Table 1. These novel approaches address challenges associated with limited clinical data by augmenting datasets, thus enhancing generalizability and mitigating overfitting in downstream machine learning tasks across various healthcare domains. In this section, we systematically examine generation methodologies among 59 reviewed studies, summarizing recent trends in model architectures, generation techniques, and integration strategies for synthetic data generation. Specifically, we aim to answer two key questions: 1) what predominant methods (e.g., prompt-based vs. specialized architectures) are currently employed for biomedical synthetic data generation; and 2) what trends and variations do exist regarding model selection (open-source vs. closed-source) and prompting strategies (zero-shot, few-shot, knowledge-infused)?
Our review identifies prompt-based generation as the predominant synthetic data generation method, employed by approximately 72.9% ($n=43$) of the analyzed studies. Prompt-based approaches [64, 31, 52] primarily rely on meticulously crafted textual instructions, leveraging fine-tuned LLMs, such as T5 and GPT variants (e.g., GPT-4) for diverse biomedical tasks. Few-shot prompting provides a small number of curated examples (typically 2-5) within the prompt to illustrate the task and guide LLMs for synthetic data generation, which count 20.3% ($n=12$). The approach uses limited curated examples to guide LLMs in generating clinically relevant synthetic data across diverse applications, including suicide symptom extractions [8], phenotype classification [45, 53, 25, 12, 65], and therapeutic dialogue generation [13, 35]. For example, Ghanadian et al. [8] use a patient’s suicidal narrative and a medication counseling dialogue as few-shot prompts to generate clinical texts for suicide symptom extraction and therapeutic dialogue generation, while Nievas et al. [32] use clinical trial inclusion and exclusion criteria paired with patient summaries as few-shot prompts to generate a synthetic patient–trial matching dataset. Zero-shot prompting involves prompting LLMs without any illustrative examples and replies solely on explicit task instructions and the pretrained knowledge encoded in LLMs [61]. This approach is suitable for simpler tasks, such as dialogue summarization [21] or generating structured medical reports [50], where extensive contextual training is less critical. For example, Barr et al. [61] used zero-shot prompting with GPT-4o to synthesize perioperative clinical tables of patient demographics, surgical parameters and outcomes, showing that the model can produce realistic synthetic data and preserve key statistical patterns based solely on qualitative instructions. Additionally, ClinGen [11] studies utilize knowledge-infused prompting, which explicitly incorporates external biomedical resources, such as clinical ontology or structured knowledge graphs. The prompts integrate the biomedical resources to improve factual accuracy, clinical relevance, and contextual richness. For example, Chuang et al. [62] construct keyword-driven prompts by appropriate ICD-9 (International Classification of Diseases, 9th Revision) codes per primary diagnosis and incorporate relevant clinical terms into prompts for GPT-3.5 and -4.
A notable trend is the model selection between open-source and closed-source LLMs, such as Llama 4 vs GPT-4. Approximately half of the reviewed studies (67.8%, $n=40$) directly utilize powerful closed-source models such as GPT-3.5 and GPT-4 through proprietary APIs. This approach prioritizes rapid development and sophisticated output quality without extensive model customization. In contrast, 39.0% ($n=23$) of prompt-based studies utilize fine-tuned open-source models (e.g., GPT-2 [68, 19, 25, 22, 24, 20], T5 [37, 8]), offering flexibility to specialize the model with biomedical domain-specific data. Fine-tuning [26, 17, 24] involves additional training of a pretrained model on a smaller, task-specific dataset, enabling the model to specialize its outputs according to specific biomedical contexts or data types. Fine-tuning open-source models demonstrates advantages in privacy-sensitive or -nuanced clinical contexts [20, 26, 22], such as assisted suicide detection [8] and cancer diagnosis [45]. For example, Lu et al. [22] fine-tuned GPT-2 on MIMIC-III [70] discharge summaries to generate synthetic clinical notes for data augmentation, effectively mitigating class imbalance, preserving patient privacy, and improving the accuracy of a downstream readmission-prediction model. However, while open-source models offer greater customization potential, they typically underperform closed-source models in general language tasks and clinical reasoning benchmarks [57, 34, 62, 8, 12, 36]. For example, Jeong et al. [36] report that the open-source model LLaMA2 7B underperforms GPT-4-base on clinical reasoning benchmarks.
Approximately 27.2% ($n=16$) of the reviewed studies explore specialized model architectures or fine-tuned pre-trained LLMs for synthetic data generation, including multimodal, multi-agent frameworks and custom-designed transformer architectures (e.g., HALO) [67, 66, 43, 55, 41, 13, 59, 27, 18]. Multimodal approaches integrate multiple data modalities, such as text, images, and audio, to create richer, more realistic synthetic datasets, such as colonoscopy image analysis [66], digital pathology interpretation [67], and emergency medical services (EMS) assistance [43]. For example, CognitiveEMS [43] prompts an LLM with emergency protocols to generate synthetic cardiac-arrest dialogues, converts them to speech, and synchronizes the audio with AR smart-glasses video frames annotated by a zero-shot vision classifier, producing a unified dataset for cardiac emergency analysis. Multi-agent frameworks [51, 29, 13] involve employing multiple interacting models or agents, each performing distinct roles, to collaboratively produce complex and realistic data, such as doctor-patient dialogues or clinical scenarios. For instance, NoteChat [13] employs a cooperative multi-agent framework with planning, role play and refinement modules to generate synthetic patient–physician dialogues that mirror real clinical encounters. Custom transformer architectures[59, 27, 18] extend the standard decoder by integrating domain-specific metadata and employing hierarchical modeling strategies, first capturing broader structural information and then refining finer details, to generate structured and realistic synthetic clinical datasets tailored to specific medical scenarios. For example, Theodorou et al. [27] demonstrate that their hierarchical autoregressive model generates privacy-preserving EHRs whose temporal and code distributions yield downstream prediction performance on par with real data. Collectively, these specialized frameworks enhance the fidelity and applicability of synthetic biomedical data for various complex health scenarios.
Table 2: Evaluation characteristics of reviewed studies: disease, downstream task, number of automated metrics (Metrics #), inclusion of intrinsic evaluation (Intrinsic+), human involvement in generation or task evaluations (Human-in-the-Loop), and LLM-based evaluation (LLM Eval). CLS, NER, IE, RE, and QA refer to downstream tasks of classification, named entity, recognition, information extraction, relation extraction, and question answering, respectively.
Study Disease Task Metrics # Intrinsic+ Human-in-the-Loop LLM Eval
[17] General QA 3 Yes No No
[18] Mental health CLS 6 Yes Yes No
[19] General NER 4 No Yes No
[20] General NER 5 No Yes No
[21] General Summarization 3 No Yes No
[22] General CLS 3 No No No
[23] General RE, QA 4 No Yes No
[24] General NER 6 Yes Yes No
[25] Febrile convulsions CLS 3 No Yes No
[26] Suicide CLS 2 No Yes No
[27] General CLS 8 No Yes No
[12] Alzheimer CLS 4 No Yes No
[28] General CLS 3 No Yes No
[29] PTSD CLS 3 No Yes No
[30] Fundus Fluorescein Angiography QA 5 No Yes No
[31] Acute Renal Failure CLS 7 No Yes No
[10] General CLS 5 No No No
[32] General CLS 8 No Yes No
[33] General NLI 5 No No No
[34] General - 1 Yes No No
[35] Psychological Therapy - 9 Yes No No
[36] General QA 3 No No No
[37] General QA 8 No Yes No
[11] General CLS, RE, NER 7 Yes No No
[38] General IE, CLS 3 No Yes No
[9] Depression Semantic Search 4 No No No
[13] General Dialogue Generation 7 Yes Yes Yes
[40] General CLS 7 No No No
[41] Gait-based Disease Pose Estimation 4 Yes No No
[42] General IE 1 Yes Yes Yes
[43] General ASR 2 No No No
[44] General NER, RE 1 No No No
[8] Suicide CLS 2 No Yes No
[45] Cancer CLS 2 No No No
[46] General Summarization 4 Yes Yes Yes
[47] General QA 1 No No No
[48] Skeleton pose CLS 2 Yes No No
[49] General - 2 Yes Yes No
[50] General Summarization 6 No No Yes
[51] General QA 4 No Yes Yes
[52] Breast cancer CLS 6 Yes No No
[39] COVID-19 CLS 1 No Yes No
[53] Autism CLS 3 No Yes No
[54] Stroke thrombolysis contraindications CLS 6 No Yes No
[55] General CLS 4 No Yes No
[56] Phenotype ontology IE 3 No No No
[57] Limb Fractures CLS 3 Yes No Yes
[58] Anesthesiology QA 5 No Yes Yes
[59] General CLS 2 No No No
[60] Mental health IE 2 No Yes No
[61] Perioperative clinical data - 1 Yes No No
[62] General CLS 3 Yes No No
[63] General QA 5 No Yes Yes
[64] General CLS 2 No No No
[65] Cancer CLS 1 No Yes No
[66] Colorectal cancer CLS, Segmentation 7 No No No
[67] Breast cancer CLS, Segmentation 4 No Yes No
[68] General NER 3 No Yes No
[69] Breast Cancer, Diabetes CLS 3 No No No
Quality Assessment and Evaluation Metrics
Ensuring quality of the synthetic data is the key to build precise models and medical applications. In this section, we assess our collected studies from the synthetic data assessment and evaluation perspective, as shown in Table 2. We aim to answer the following three questions: 1) What metrics and evaluation approaches are commonly used for assessing synthetic data quality? 2) What are major challenges in model evaluations using synthetic data across biomedical diseases and tasks; and 3) What are the emerging trends in synthetic data quality control for biomedical research? To answer those questions, we systematically examined the collected 59 studies by disease topics, types of evaluation metrics (automated and human approaches), downstream biomedical tasks, and emerging strategies (e.g., LLM-based and intrinsic evaluations). We aim to highlight current practices, identify persistent challenges, and discuss notable shifts toward more robust and clinically meaningful data and model assessments in the biomedical synthetic data generation.
We examined two major evaluation metrics across diverse biomedical applications, intrinsic and extrinsic metrics, which are the essential measurements to select synthetic data and computational models. Intrinsic metrics (e.g., perplexity [42] or similarity measures such as KL Divergence [57]) assess properties of the synthetic data independently from downstream tasks, while extrinsic metrics rely on downstream tasks, such as classification [18]. As shown in Table 2, we can observe that the intrinsic metrics are not widely adopted and only count 27.1% ($n=16$) of the studies. For example, Study [11] and Study [46] employ intrinsic metrics to directly compare the statistical distribution of real and synthetic datasets, while most works, such as Studies [22, 25], and [56], rely primarily on extrinsic evaluation through classification (CLS) or information extraction (IE) tasks. Commonly used metrics include accuracy [26], F1 score [25], BLEU [30], and other task-specific measurements, such as ROUGE [63] and AUCROC [22]. The studies cover heterogeneous downstream tasks, such as or de-identification for privacy protection [10], phenotype prediction [56], disease entity extraction [42], and clinical note summarization [46, 21]. Privacy protection assessment was reported in 20.3% ($n=12$) of the studies, primarily through membership and attribute inference attacks [10], adversary test [27], and memorization evaluation [18]. Those studies cover multiple medical domains, with most studies focusing on general biomedical applications (n=33) and others targeting 25 specific medical conditions and health-related topics, such as Alzheimer’s [12], breast cancer [52], and depression [9].
Beyond automated metrics, human evaluation has played a selective yet important role in the assessment of synthetic biomedical data or its utility in different tasks (e.g., phenotype inference or clinical note summary), particularly for tasks where clinical relevance or nuanced interpretation is required. Human evaluation [32, 12, 13] typically involves domain experts to determine if synthetic data are meaningful, such as clinicians or biomedical researchers, who assess the synthetic data for factors like clinical validity. Human evaluation was included in a minority of studies (44.06%). Those studies cover tasks such as clinical entity recognition [19], clinical note summarization [21, 46], or question answering [37], where expert judgment adds valuable context beyond quantitative scores. More recently, a few studies [13, 42, 46, 50] (13.55% of our studies) explore using large language models (LLMs) as automated evaluators and leverage their ability to perform nuanced judgments, which approximates human assessment. Common LLM-based evaluation metrics employed in these studies include factual consistency [50], clinical correctness [42], and error identification [51], often measured through prompt-based scoring or rubric-guided assessments conducted by the LLM itself. For example, a recent study [50] prompts GPT-4 if the context supports / contradicts the response, or if there’s not enough information as a hallucination evaluation, whereas another study [42] uses a rubric-guided prompt, requesting GPT-4 to score each response on a four-level scale, Unacceptable, Poor, Satisfactory, and Excellent. However, LLM-based evaluations remain an emerging approach and have not yet replaced the need for statistical metrics or domain expert involvement, especially in clinically sensitive or complex scenarios [71].
Despite substantial progress, several challenges persist in the quality control of biomedical synthetic data. A key issue is the lack of standardized evaluation frameworks, resulting in wide variability in both the selection and reporting of metrics across studies and disease domains [11]. This heterogeneity complicates direct comparison between methods and limits the transferability of findings across medical applications and diseases. Additionally, Table 2 shows that studies still underutilize intrinsic evaluation, focusing instead on extrinsic, task-specific benchmarks that may not fully capture utility or limitations of synthetic data. Emerging trends suggest a gradual shift toward more comprehensive assessments, including extrinsic and intrinsic evaluation integrations, LLM-based metrics adoptions, and human professional engagements.
Discussion
Our scoping review has presented the current developments and applications of biomedical synthetic data generation, which is being facilitated by large language models (LLMs). This section, we highlight several insightful takeaways of emerging methodological and evaluative trends and identify promising areas for future biomedical research and applications using synthetic data generation.
This review shows a clear shift towards leveraging synthetic data to overcome biomedical research challenges, such as data scarcity, privacy restrictions, and diverse clinical conditions. While synthetic data has been used effectively across varied clinical applications (e.g., phenotype ontology extraction [56] and cancer diagnosis [66, 67, 45]), the scale still vary across data types and disease fields, reflecting the nuanced capabilities and limitations of current LLMs. Clinical narratives and unstructured texts remain the most common data source, due primarily to their widespread availability and the inherent LLMs strengths in natural language processing tasks. We can also find an emerging trend toward multimodal and structured synthetic EHR, which indicates increasing expansions across more clinical settings, such as imaging, audio, and wearable sensors [41]. However, the advances indicate potential concerns in methodological developments. Our review shows that current studies are significant depending on prompting closing-sourced LLMs, like GPT-4. While few-shot and zero-shot approaches may achieve some successful cases [29, 44, 31], prompting closing-sourced LLMs (e.g., GPT-4) can limit transparency, reproducibility, and the ability to customize models for specialized clinical contexts. Additionally, access to proprietary models may be restricted by cost, availability, or data governance policies. For example, MIMIC-IV (Medical Information Mart for Intensive Care) [72] does not allow for uploading data to those closing-sourced LLMs. Future work should address these concerns by promoting the development and benchmarking of open-source models and by exploring alternative methods that reduce dependence on proprietary LLMs.
Our review indicates future research in biomedical synthetic data generation may evolve along three dimensions. There is a clear trend towards generating multimodal synthetic datasets to simulate more real-world and complex clinical scenarios, such as radiology [57, 65], pathology [67], and emergency care [34]. Future synthetic data generation may depend on more precise LLM-based systems by knowledge-guided generations [11], multi-agent frameworks [13], and human-LLM collaborations [39], which leverage clinical knowledge to enhance factuality and specificity of synthetic data qualities. In terms of evaluation, more standardized, multi-dimensional, and human-centered assessments are critical. Future studies on biomedical synthetic data generation may consider a mixture of intrinsic and extrinsic metrics, human-in-the-loop approaches, and task-agnostic validations to ensure fidelity, utility, and privacy of synthetic data. We envision the directions will essential to advance biomedical research and applications by the synthetic data generations and LLMs.
Limitations
This scoping review has several limitations that should be acknowledged to appropriately interpret our findings and analysis. First, the literature search was limited to specific databases and publication venues, potentially missing relevant studies published in less prominent journals or literature sources. Second, our review primarily included studies written in English, potentially overlooking important findings published in other languages. Additionally, the rapid pace of research in LLM-driven synthetic data generation means that emerging studies and recent methodological advances might not be captured fully in this review. Lastly, given the inherent heterogeneity in study designs, evaluation metrics, and application domains across the reviewed literature, direct comparison and generalization of findings are challenging, underscoring the need for more standardized reporting and evaluation frameworks in future research.
Methods
Data Source and Search.
We conduct searches between April and May 2025 following the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) guidelines [73] and the PRISMA Extension for Scoping Reviews (PRISMA-ScR) [74] (Figure 2). Recognizing the rapid developments of Large Language Models (LLMs) in the biomedical field, our review encompassed peer-reviewed articles between January 1, 2020, to April 5, 2025 (inclusive) from multiple database sources, including ACM Digital Library, PubMed, Web of Science, and Google Scholar. Supplemental articles were gathered by reviewing article bibliographies and by soliciting suggestions from XH, HW, IH, and ZH. Our search strategy utilized a logical combination of keywords: “Language Model” AND “synthetic data” AND “medical OR health OR clinical”. We primarily limited the search to titles and abstracts, expanding to full texts where this function was unavailable. For Google Scholar, we conducted a keyword search, sorted by relevance, and selected the first 100 studies. The search strategies for each database were initially formulated in the early stages of the study and refined through team discussions and preliminary analysis of the results.
Figure 2: PRISMA-ScR flow diagram.
Study Selection.
We followed inclusion and exclusion criteria to narrow the candidate articles. Titles and abstracts were initially screened against predefined eligibility criteria by at least two independent reviewers, excluding records that clearly did not meet inclusion criteria and retaining ambiguous ones for full-text screening. Our inclusion criteria include: 1) an article conducts biomedical research; 2) an article uses Large Language Models (LLMs) for synthetic generation of biomedical data; and 3) an article has undergone peer review and been published in a journal or in conference proceedings within the range of 01/01/2020 and 04/05/2025. Our exclusion criteria include: 1) an article is a survey, literature review, news article, editorial, letter, opinion, study protocol, and comment; 2) an article is not in English; 3) articles do not study biomedical large language models; and 4) an article is not peer-reviewed and published in a journal or conference proceedings. Three undergraduate students from the XH lab assisted with review of candidate articles with the HS and XH, who divided the candidate articles into relatively equal numbers and assigned them to reviewers. Reviewers examined titles, abstracts, full-text, and other article metadata, recommending inclusion or exclusion based on the defined criteria. Each article was reviewed by at least two individuals, and disagreements were resolved after further review by XH and HW.
Data Extraction.
Selected articles were mainly processed by two individuals (HR and XH), who are experts in natural language processing (NLP) and biomedical informatics. They independently reviewed the full text of each selected article and extracted key metadata. The extracted data focused on three main aspects: 1) Synthetic data types and applications, including data characteristics , such as language and domain; 2) Synthetic data generation methods, presenting generation pipelines (models used to generate synthetic data) and their data modalities (e.g., text, image, audio); 3) Quality assessment and evaluation metrics summarizing medical applications, evaluation approaches, and evaluation settings. Discrepancies in data extraction were resolved through discussion between HR and XH, with further arbitration by XH, HW, and IH when necessary. All extraction processes were documented for replication purposes.
Acknowledgments
WR, WL, and WH were partially supported by the National Science Foundation under Award IIS-2245920 and IIS-2440381, and the work was partially supported by National Cancer Institute Award R01CA258193. ZH was supported by the Agency for Healthcare Research and Quality Award R21HS029969, the National Library of Medicine Award R21LM013911, the National Institute of Mental Health Award R21MH137736, and the National Institute on Aging Award R01AG064529. The authors would like to thank three undergraduate students, Sampoorna Poria, Sashank Reddy Karri, and Precious T Jones, for assisting paper collection and filtering at the initial stage.
Author contributions statement
H.R. and W.L. led the literature search, data extraction, and formal analysis. H.R. and W.L. made equal contributions to this study. H.W. contributed to initial data collection, methodology development, and manuscript preparation. W.L. joined the process at a later stage. IC.H. and Z.H. provided domain expertise, supervised the review process, and revised the manuscript. X.H. conceived and designed the study, coordinated the project, and contributed to writing, editing, and finalizing the manuscript. All authors reviewed, edited, and approved the final manuscript.
Additional information
Competing interests
The authors declare no competing interests.
Data availability
Not applicable.